OpenStudy (anonymous):

How do I find the volume if I don't have the height???? Kent has two similar cylindrical pipes, Pipe A and Pipe B. The radius of Pipe A is 6 cm, and the radius of Pipe B is 2 cm. What is the ratio of the volume of Pipe A to the volume of Pipe B?

OpenStudy (lgbasallote):

you dont need to find the volume...you just need the ratio

OpenStudy (lgbasallote):

volume of pipe A is \[V_A = \pi r^2 h\] \[\implies V_A = \pi (6)^2 h\] \[\implies V_A = 36\pi h\] agree?

OpenStudy (anonymous):

Yes. The other will be v=pi(2)^2h which turns out to be v=4pih

OpenStudy (lgbasallote):

right. so the ratio of pipe a to pipe b is \[\huge \frac{V_A}{V_B} \implies \frac{36\pi h}{4\pi h}\] simplify that to get the ratio ;)

OpenStudy (anonymous):

The h and the pi are going to cancel out then?

OpenStudy (lgbasallote):

yup

OpenStudy (anonymous):

So the answer is 9:1

OpenStudy (lgbasallote):

correct you are

jimthompson5910 (jim_thompson5910):

hmm....

jimthompson5910 (jim_thompson5910):

if the solids are similar, then wouldn't the second height be 3 times larger than the first?

jimthompson5910 (jim_thompson5910):

since every corresponding dimension is in proportion with each other

jimthompson5910 (jim_thompson5910):

also, I'm thinking that the volume ratio should be 27:1 since 3^3 = 27

OpenStudy (lgbasallote):

yes...i was wondering the same thing...it should be \[\frac{216}{8}\] wonder where it went wrong...the h maybe @jim_thompson5910 ?

jimthompson5910 (jim_thompson5910):

oops nvm, miscalculated

OpenStudy (lgbasallote):

216/4 gives 54..

jimthompson5910 (jim_thompson5910):

yeah my bad

jimthompson5910 (jim_thompson5910):

you're right, 216/8 = 27

OpenStudy (lgbasallote):

hmm...the ratio of volume is supposed to be the cube of the ratio of radii

OpenStudy (lgbasallote):

so how would it look like algebraically?

OpenStudy (lgbasallote):

maybe \[\huge V_A = \pi(6)^2 (3h)\] and \[\huge V_B = \pi (2)^2(h)\] ??

jimthompson5910 (jim_thompson5910):

The larger radius is 3 times the smaller radius (since 6/2 = 3) So each dimension of the larger cylinder is triple its corresponding dimension so if r1 and h1 are the radius and height of the smaller cylinder and r2 and h2 are the radius and height of the larger cylinder, then.... r2 = 3*r1 h2 = 3*h1 which makes the volumes to be V1 = pi*(r1)^2*h1 V2 = pi*(r2)^2*h2 and the ratio of the volumes to be (V2)/(V1) = (pi*(r2)^2*h2)/(pi*(r1)^2*h1) but r2 = 3*r1 and h2 = 3*h1, which means we can say (V2)/(V1) = (pi*(3*r1)^2*(3*h1))/(pi*(r1)^2*h1) (V2)/(V1) = ((3*r1)^2*(3*h1))/((r1)^2*h1) (V2)/(V1) = ((3*r1)^2*(3))/((r1)^2) (V2)/(V1) = (9*(r1)^2*(3))/((r1)^2) (V2)/(V1) = (9*3)/(1) (V2)/(V1) = 27

jimthompson5910 (jim_thompson5910):

So the ratio of the volumes is 27:1

jimthompson5910 (jim_thompson5910):

yes, what you just wrote is exactly what I'm thinking

OpenStudy (lgbasallote):

ahh yes...that's how it should be @tiffybabyy sorry for the mistake

jimthompson5910 (jim_thompson5910):

your prev answer would work if the two heights were the same...but that would mean that the two solids aren't similar

jimthompson5910 (jim_thompson5910):

hopefully I didn't needlessly complicate things for you tiffybabyy lol

OpenStudy (anonymous):

Okay no I'm following.

jimthompson5910 (jim_thompson5910):

alright, let me know if you get lost anywhere